### v1

# mmCooper: A Multi-agent Multi-stage Communication-efficient and Collaboration-robust Cooperative Perception Framework

Abstract
Collaborative perception significantly enhances individual vehicle perception performance through the exchange of sensory information among agents. However, real-world deployment faces challenges due to bandwidth constraints and inevitable calibration errors during information exchange. To address these issues, we propose mmCooper, a novel multi-agent, multi-stage, communication-efficient, and collaboration-robust cooperative perception framework. Our framework leverages a multi-stage collaboration strategy that dynamically and adaptively balances intermediate- and late-stage information to share among agents, enhancing perceptual performance while maintaining communication efficiency. To support robust collaboration despite potential misalignments and calibration errors, our framework captures multi-scale contextual information for robust fusion in the intermediate stage and calibrates the received detection results to improve accuracy in the late stage. We validate the effectiveness of mmCooper through extensive experiments on real-world and simulated datasets. The results demonstrate the superiority of our proposed framework and the effectiveness of each component.
1 Introduction
With the development of autonomous driving systems and Internet of Vehicles (IoVs) technologies, perception devices such as cameras and LiDARs have been widely deployed. Although perception, as a critical module in autonomous vehicles, has witnessed rapid advancements driven by deep learning-based algorithms, the traditional single-vehicle perception paradigms [19, 24, 45] cannot meet the safety and reliability requirements of autonomous vehicles due to unavoidable factors such as object occlusion and limitations in detection range [53, 1, 12, 4, 22].
Figure 1: Comparison of different fusion approaches.
Benefiting from infrastructure improvements and the advancement of communication technologies like V2X [21, 20], autonomous vehicles can achieve multi-vehicle collaborative perception by sharing perception information [39, 3], which significantly enhances the performance of perception systems. Although cooperative perception has made progress recently, its practical deployment still faces challenges such as communication resource constraints [11, 6], localization errors [25, 26] and low information fusion efficiency [48, 38].
In cooperative perception systems, the effectiveness of fusion strategies significantly affects both perception performance and communication efficiency. As illustrated in Figure 1 (a)-(c), the fusion strategies can be divided into three categories: early, intermediate, and late fusion methods, illustrated. Specifically, early fusion methods [5, 53] aggregate raw sensor observations from all agents, providing abundant information that can significantly enhance perception performance. However, transmitting raw sensor observations requires substantial bandwidth, making this approach costly for communication. Intermediate fusion methods [11, 6, 48, 47, 46] address this issue by sharing encoded features rather than raw data to reduce communication overhead. While effective, feature compression inevitably introduces information loss, which can compromise fusion accuracy. Late fusion methods [30, 32] on the other hand, minimize communication demands by sharing lightweight perception results among agents. However, this approach is highly sensitive to environmental noise, as even minor disturbances can degrade collaborative quality. Overall, these above single-stage cooperative perception approaches face challenges in managing the trade-off between perception accuracy and communication efficiency, raising a compelling question: can we improve the perception performance while preserving communication efficiency by strategically sharing information across multiple stages to harness the strengths of each fusion stage?
In addition, the data exchanged by agents inevitably experiences misalignment and contains calibration errors during the synchronization process due to communication delays and pose noise [13, 33]. These calibration errors can produce misleading features and proposals in subsequent steps, thereby affecting the performance of the perception system. While existing intermediate fusion methods attempt to address the calibration errors through position-wise feature fusion [18, 16, 23], they often overlook the critical role of neighboring region information for enhancing collaborative robustness. Meanwhile, late fusion methods [2, 28, 34] simply merge bounding boxes from collaborative agents, failing to leverage information-rich intermediate features from the ego agent, which could help refine and calibrate these bounding boxes for higher accuracy.
To this end, we propose mmCooper, a novel multi-agent, multi-stage, communication-efficient, and collaboration-robust cooperative perception framework. The proposed mmCooper framework facilitates effective information aggregation across fusion stages and enhances communication efficiency through carefully designed core components. Specifically, we design an adaptive multi-stage data fusion mechanism that facilitates multi-agent cooperation by dynamically fusing information at both intermediate and late stages, as shown in Figure 1. Guided by a confidence-based filtering gaile strategy, this mechanism automatically assesses and determines the data volume to be processed at intermediate and late stages. High-confidence bounding boxes are shared directly, while intermediate features are selectively shared instead of bounding boxes for regions requiring further contextual reference, and low-confidence areas are excluded to prevent propagating misleading information. This dual-stage fusion approach enhances perception performance while maintaining communication efficiency.
Besides, to further address potential misalignment and calibration noise among agents, we incorporate specific designs in the proposed mmCooper framework for both intermediate and late stages. In the intermediate stage, we introduce a Multi-scale Offset-aware Fusion Module that not only fuses data from target locations related to cooperative agents’ views but also from nearby regions, adding redundancy to mitigate calibration noise. For the late stage, we design a Bounding Box Filtering & Calibration Module that uses the information-rich intermediate-stage features from the ego agent to filter inaccurate bounding boxes received from other agents and refine the remaining noisy ones. By integrating these corrected and ego-generated bounding boxes, our mmCooper framework is able to achieve high perception accuracy and robustness. We evaluated the performance of our mmCooper framework on both real and virtual datasets, including DAIR-V2X [49], OPV2V [43] and V2XSet[42]. Experimental results demonstrate that mmCooper outperforms previous methods across various settings, confirming its effectiveness as a communication-efficient and robust multi-stage cooperative perception framework. To summarize, our main contributions are summarized as follows:
  • We propose mmCooper, a multi-agent multi-stage communication-efficient collaboration-robust cooperative perception framework. To our knowledge, this is the first framework to address the tradeoff between limited communication bandwidth and desired perception performance by sharing both intermediate and late-stage information for multi-agent collaborative perception.
  • To address the potential misalignment and calibration error among agents, we incorporate a Multi-scale Offset-aware Fusion Module to integrate spatially adjacent contextual information at different feature scales in the intermediate stage and a Bounding Box Filtering & Calibration Module to filter and improve the bounding boxes from other agents in the late stage, guided by the information-rich intermediate features.
  • Extensive experiments show excellent performance on both real-world and simulated datasets, with a 7.29%/1.31%/2.09% improvement in AP@0.7 on OPV2V, DAIR-V2X and V2XSet datasets over the second-best state-of-the-art methods. Additionally, our approach requires only 1/9153, 1/156 and 1/18305 of the communication volume used by comparable methods.
2 Related Work
Figure 2: Overview of mmCooper framework: a cooperative perception system with adaptive multi-stage fusion.
2.1 Multi-Agent Communication
Communication strategies in multi-agent systems have been widely studied [31]. Early works [36, 29, 17] often used predefined protocols or heuristic methods to determine how agents communicate with each other. However, these methods are difficult to generalize to complex tasks. Recently, several learning-based communication strategies have been proposed to adapt to diverse scenario applications. For instance, Vain [10] utilizes attention neural structures to specify the information that needs to be shared in agent interactions. ATOC [14] introduces recurrent units to decide who agents communicate with by receiving local observations and action intentions from other agents. TarMAC [7] designs an architecture oriented towards reinforcement learning, learning to communicate from task-specific rewards. CommNet [35] learns continuous communication in multi-agent systems. Most of these previous works have considered decision-making tasks and have employed reinforcement learning due to the lack of explicit supervision. Different from these works, our paper focuses on LiDAR-based 3D perception tasks in complex autonomous driving scenarios and achieves efficient communication between agents using a multi-stage communication strategy.
2.2 Collaborative Perception
Collaborative perception addresses the limitations of single-agent perception by aggregating perception information from surrounding agents to adapt to complex and dynamic autonomous driving environments. Current work is categorized into early, intermediate, and late fusion based on collaboration stages. Early fusion [5, 51, 53] shares raw point cloud data among agents, achieving high perception performance but consuming significant bandwidth resources. EMP [53] achieved scalability and adaptability on highly fluctuating open roads by dynamically partitioning the point cloud range shared by agents. Late fusion [30, 44, 32] shares lightweight proposals between agents but performs poorly and is sensitive to noise. VIPS [30] implemented efficient graph-structured matching to enable object-level fusion. Intermediate fusion [23, 11, 42, 50, 18] shared features among agents, yet complete Intermediate features still incur high bandwidth consumption. Recent work focused on novel communication mechanisms to filter Intermediate features and reduce communication volume. for instance, Where2comm [11] employed a confidence-based scheme to guide agents in focusing on sharing spatially critical information. ERMVP [50] adopted a hierarchical feature sampling strategy to reduce communication overhead while leveraging sparse consensus features to mitigate localization errors. Different from these works, our framework conducts multi-stage fusion instead of single-stage which balances the perception performance and communication overhead. As far as we know, only a few works have considered multi-stage settings. Zhang et al. [52] proposes a defense method that combines raw data and intermediate features to fuse redundant information, thereby reducing the impact of spurious data. Zhao et al. [54] simultaneously considers early raw data and late-stage detection results to address malicious agents in collaborative environments. However, both works focus on security issues related to collaborative perception instead of improving sensing performance. Xie et al. [40] applies reinforcement learning for dynamic partitioning of early, intermediate, and late-stage information. However, this work merely divides the data without integrating information across stages. In contrast to these works, this paper introduces a collaborative perception framework that aims to improve sensing performance by conducting multi-stage data fusion.
3 Methodology
3.1 Overview
This section introduces our proposed mmCoop framework, as illustrated in Figure 2, which comprises four main components: the Observation Module, the Broadcasting Module, the Intermediate-stage Fusion, and the Late-stage Fusion. The Observation Module (Sec.3.2) serves as the foundation by extracting essential features from the ego vehicle’s LiDAR-based point cloud data. The Broadcasting Module (Sec.3.3) includes a Detection Decoder to generate coarse bounding boxes from intermediate features, along with a Confidence-based Filter Generation Module to produce filters that dynamically determine the data volume shared at the intermediate and late stages. These generated filters enable adaptive data sharing: high-confidence bounding boxes are shared directly, intermediate features are selectively transmitted for areas needing additional context, and low-confidence areas are excluded, balancing perception accuracy with communication efficiency. The Intermediate-stage Fusion (Sec.3.4) contains the Multi-scale Offset-aware Fusion Module, which uses cross-attention to integrate features from both target and nearby regions for robust feature fusion. In the Late-stage Fusion (Sec.3.5), a Bounding Box Filtering and Calibration Module filters inaccurate bounding boxes from other agents and calibrates the remaining ones using the ego agent’s intermediate features. The Late-stage Fusion contains the Bounding Box Filtering and Calibration Module to filter out inaccurate bounding boxes received from other agents and refine the remaining noisy ones by leveraging the information-rich intermediate features from the ego agent. Then, the calibrated bounding boxes from all other agents will be combined with the bounding boxes generated by the ego agent to produce the final prediction.
3.2 Observation Module
In a multi-agent cooperative perception scenario, consider the collaboration of $n$ agents, represented by the agent set $N=\left\{1,...,n\right\}$. One vehicle serves as the ego vehicle and the remaining $n-1$ agents act as collaborators. The ego vehicle aggregates information from the collaborators to achieve cooperative perception. Considering the i-th agent as the ego vehicle, for the $j$-th collaborator, where $i,j\in N$, the point cloud observation is given as $X_{j}$. The ego vehicle first broadcasts its basic information to the collaborators, such as coordinates and heading angle. The collaborators utilize this basic information to project their point cloud observations into the ego vehicle’s coordinate system, with the projected point cloud observation represented as $x_{j}^{{}^{\prime}}=T_{j->i}\cdot x_{j}$, where $T_{j->i}$ denotes the transformation matrix from agent j to agent i. Subsequently, each agent extracts features from its point cloud observation through a shared encoder to obtain Bird’s Eye View (BEV) features, represented as $F_{bev}^{k}=\psi_{encoder}\left(x_{k}\right)\in\mathbb{R}^{C\times H\times W}$, where $\psi_{encoder}$ denotes the shared PointPillar encoder [15], and $k\in N$, and $C$, $H$, and $W$ correspond to the channel, height, and width of the feature map, respectively.
3.3 Broadcasting Module
Figure 3: The Confidence-based Filtering Generation Module
For the locally encoded BEV features, prior work has mainly focused on either directly sharing intermediate-stage features or decoding them to obtain late-stage bounding boxes, which are then shared. However, collaboration schemes based on a single stage often struggle to effectively balance communication bandwidth and perception performance. To effectively leverage the advantages of multi-stage data, we introduce a multi-stage collaboration framework that simultaneously balances bandwidth and accuracy by transferring filtered features and proposals between agents. As shown in Figure 3, our Broadcasting Module filters the intermediate features and initially generates coarse bounding boxes, which are then transmitted to other agents.
Confidence-based Filter Generation Module. The Confidence-based Filter Generation Module, as a key component of the Broadcasting Module, serves two main purposes: 1) dynamically and adaptively determining the allocation of features and coarse bounding boxes; 2) ensuring that the discrete operations involved in obtaining the filter can still propagate gradients properly. To filter spatial features and bounding boxes, we input the encoded features into the Confidence Generator to generate spatial confidence maps for both the features and proposals. The confidence maps represent the potential regions of objects and are denoted as:
$C_{f}^{i},C_{b}^{i}=\phi_{conf}\left(F_{bev}^{i}\right),$
where $\phi_{conf}\left(\cdot\right)$ refers to the Multi-stage Confidence Generator, and $C_{f}^{i},C_{b}^{i}$ represents the confidence maps for the features and proposals, respectively. A Gaussian Filter is then applied to the spatial confidence maps, helping to filter out anomalous areas and incorporate contextual effects, thereby enhancing the robustness of critical region selection. To eliminate irrelevant background information, the top $m\%$ highest-confidence positions are selected from each confidence map to generate the spatial filter $M_{f/b,top}^{i}$. Then we add Gumbel noise $g_{f/b}$ to the confidence maps to introduce a certain level of randomness, expressed as:
$g_{f/b}=-log\left(-logu\right),u\sim Uniform\left(0,1\right).$
The introduction of Gumbel noise can enhance the exploratory nature of the model and improve its generalization capability. For the same positions in the confidence maps from two stages, we retain only the parts with relatively high confidence scores. The confidence score indicates the potential location of an object. When a high-dimensional bounding box exhibits high confidence in representing the object, the agent will share the bounding box. Otherwise, intermediate-stage features are required to convey the information at that location. In addition, to ensure the continuity of the dynamic allocation process for the intermediate-stage features and coarse bounding boxes. we employ the Gumbel Softmax method to obtain approximate samples from a discrete distribution. Utilizing the straight-through estimator to acquire discrete distributions while preserving standard forward propagation. The specific process is represented as follows:
$M_{f/b,g}^{i}=f_{max}\left(\frac{exp\left((logC_{f/b}+g_{f})/\tau\right)}{{{\textstyle\sum_{s}}exp\left((logC_{s}+g_{s})/\tau\right)}}\right),$
where $\tau$ denotes the temperature parameter of the distribution $(\tau>0)$, and $s\in{\left\{f,b\right\}}$. Based on the above process, our final filter $\mathcal{M}_{f/b}^{i}$ is represented as:
$\mathcal{M}_{f/b}^{i}=M_{f/b,g}^{i}\odot M_{f/b,top}^{i},$
3.4 Intermediate-stage Fusion
Figure 4: (a) The Multi-scale Offset-aware Fusion Module. (b) The Multi-scale Offset-aware Attention Module.
The data shared between agents inevitably contains calibration errors, which cause misalignment of feature maps. These errors may lead to the ego vehicle misjudging the positions of surrounding objects, thereby reducing detection performance. To address these issues, we implement a novel Multi-scale Offset-aware Fusion (MOF) Module in the intermediate-stage Module. By aggregating information from surrounding positions, the per-position feature fusion process introduces redundancy, allowing it to effectively tolerate calibration errors.
Multi-scale Offset-aware Fusion. The overall pipeline of this module is shown in Figure 4(a). First, the ego features and the received collaborator features are encoded into three scales to capture target information of varying sizes. Then, our multi-scale offset-aware attention is applied for feature fusion. Upsampling is performed on different scales to obtain fused features of the same size. The details of the Multi-scale Offset-aware Attention Module are shown in Figure 4(b). We encode the ego features into queries via a convolutional layer, and employ the spatially continuous features at the query positions of each collaborator’s features as keys and values in the cross-attention mechanism. Define the specified location as $(i,j)$ on the feature map, with a neighborhood size of $s\times s$. The output representation of feature fusion at location $c$ is expressed as:
$\mathcal{F}_{i,c}=CrossAttn(\sigma(F_{bev,c}^{i}),F_{bev,c}^{j,s\times s},F_{bev,c}^{j,s\times s}),$
[…]
3.5 Late-stage Fusion
Due to the limited perception capabilities of individual agents, they often generate imprecise or even incorrect bounding boxes. Existing late-stage fusion methods focus only on the fusion strategy at the current stage. By utilizing data sharing across multiple stages, we are able to fully utilize the abundant intermediate fused features to calibrate the received bounding boxes.
Figure 5: (a) The BBox Filtering & Calibration Module. (b) The Deformable Bounding Box Attention Module.
BBox Filtering & Calibration Module. In this module, we propose a Feature-guided Bounding Boxes Calibration Transformer (FBCFormer) to calibrate the bounding boxes transmitted by collaborators and filter out erroneous ones. As depicted in Figure 5(a), we use the received bounding boxes from collaborators and the intermediate-stage fused features as inputs to the Deformable Bounding Boxes Attention (DBA) Module. The remaining components follow the vanilla transformer setup [37]. We apply score and offset projections to the output of the FBCFormer to learn the scores and offsets for each bounding box, thereby removing incorrect bounding boxes and calibrating the coarse ones.
Figure 5(b) provides a detailed illustration of our DBA module. We first encode the bounding boxes received from collaborators. By applying our bounding boxes encoder to the bounding box coordinates, size, and orientation information, we effectively extract the bounding box feature. We use a deformable cross-attention module to focus on keypoints information, which yields a robust representation of bounding box errors. Specifically, the bounding box features serve as the initial query embedding $q$, and the bounding box coordinates act as reference points. A linear layer encodes each bounding box’s coordinates into position embeddings. Subsequently, the linear layer learns the 2D spatial offsets for each reference point, extracting keypoints information at $q+\bigtriangleup q$. The cross-attention layer aggregates keypoint information from multiple collaborator features to enhance the representation for each query embedding:
$DBA(q)=\sum_{\alpha=1}^{A}W_{\alpha}[\sum_{n=1}^{N}\sum_{m=1}^{M}\omega(W_{\beta}F_{b}(q))F_{j}(q+\bigtriangleup q_{m})],$
where $A$ is the number of attention heads, $W_{\alpha/\beta}$ are learnable parameters, $m$ represents the number of keypoints, and $\omega$ denotes the softmax operation. Finally, two linear layers are learned to determine the offsets and scores for each bounding box. The offsets are used to adjust the position, size, and angle of each bounding box, while the scores are filtered using a threshold selection function $f_{sel}(\cdot)$ to discard lower-scoring bounding boxes. To prevent learning unreasonable offsets and scores, we control the output range using the tanh activation function. Additionally, to train our FBCFormer component, we compute the offset loss $\mathcal{L}_{off}$ using the smooth absolute error loss between the detected bounding boxes and the corresponding ground truth bounding boxes with the same identifier. The score loss $\mathcal{L}_{score}$ is calculated using focal loss.
3.6 Loss Functions
We use two detection decoders to perform regression and classification on the fused features. The regression output represents the position, size, and yaw angle of each predefined box, expressed as $O_{reg}=f_{dec}^{r}(\mathcal{F}_{i})\in\mathbb{R}^{7\times H\times W}$. The classification output indicates the confidence score for each predefined box, represented as $O_{cls}=f_{dec}^{c}(\mathcal{F}_{i})\in\mathbb{R}^{2\times H\times W}$. To optimize the proposed system, we use smooth absolute error loss for regression and focal loss for classification, denoted as $\mathcal{L}_{reg}$ and $\mathcal{L}_{cls}$, respectively. Combining the above losses, our total loss is represented as:
$\mathcal{L}=\mathcal{L}_{reg}+\mathcal{L}_{cls}+\mathcal{L}_{off}+\mathcal{L}_{score}.$
Models Venue OPV2V DAIR-V2X V2XSet
Comm AP@ Comm AP@ Comm AP@
$(log2)(\downarrow)$ 0.7/0.5 ()($\uparrow$)( ↑ ) $(log2)(\downarrow)$ 0.7/0.5 ()($\uparrow$)( ↑ ) $(log2)(\downarrow)$ 0.7/0.5 ()($\uparrow$)( ↑ )
No Fusion* [15] CVPR 2019 - 48.66/68.71 - 43.57/50.03 - 40.20/60.60
Late Fusion* [15] CVPR 2019 10.36 59.48/79.62 9.08 34.47/51.14 10.26 30.75/54.92
Intermediate Fusion* [15] CVPR 2019 23.22 70.82/ 88.41 22.62 39.38/56.22 23.22 59.38/83.18
When2com [23] CVPR 2020 23.22 57.55/74.11 22.62 33.68/48.20 23.22 41.85/67.41
DisoNet [18] NIPS 2021 23.22 68.64/84.72 22.62 40.69/52.67 23.22 54.11/79.82
Where2comm [11] NIPS 2022 23.22 69.73/85.16 22.62 43.71/59.52 23.22 63.77/83.17
V2X-ViT [42] ECCV 2022 23.22 70.06/84.65 22.62 40.43/53.08 23.22 61.49/83.63
ERMVP [50] CVPR 2024 22.46 64.60/85.81 21.98 46.96/64.21 18.46 58.47/81.14
Ours - 10.09 78.11/ 88.93 15.33 48.27/ 65.12 9.06 65.86/84.40
Table 1: Collaborative perception performance on the OPV2V, DAIR-V2X and V2XSet dataset with a time delay of 100 $ms$, localization errors of 0.2 $m$, and heading errors of 0.2° using Average Percision(AP)@0.7/0.5 as metrics. $Comm$ denotes the communication volume. Bold numbers indicate the best results, while underlined numbers represent the second-best results. * represents a variation of PointPillar [15].
4 Experiments
4.1 Datasets and experimental settings
Datasets. We conduct extensive validation of our proposed mmCooper on three benchmark datasets, namely OPV2V [43], DAIR-V2X [49] and V2XSet[42]. OPV2V is the first large-scale public V2V cooperative perception simulation dataset, collected using the autonomous driving simulator CARLA [8] and the cooperative driving simulation framework OpenCDA [41]. This dataset includes 6 types of road scenarios and 9 different cities. It comprises 73 different scenes, 11,464 frames of point clouds and RGB images, and over 230,000 annotated 3D detection boxes. The training, validation, and test sets are divided into 6,374, 1,980, and 2,170 frames, respectively. DAIR-V2X is a comprehensive large-scale real-world 3D object detection dataset designed for vehicle-road cooperative autonomous driving research. This dataset includes 71,254 frames of point cloud and image data, with the training, validation, and test sets split in a 5:2:3 ratio. The V2XSet dataset is a large-scale synthetic dataset designed for V2X perception. It consists of 73 scenarios collected using CARLA and OpenCDA. It contains a total of 11,447 frames, with each frame involving up to five agents. The dataset is divided into training, validation, and test sets, consisting of 6,694, 1,920, and 2,833 frames, respectively.
Evaluation metrics. To evaluate the 3D object detection performance of the baseline and the proposed framework, we use the average precision (AP) [9] at Intersection-over-Union (IoU) thresholds of 0.5 and 0.7 as the evaluation metric. The communication volume between agents is expressed by the following equation:
$\textbf{V}_{comm}=\log_{2}{((\left|M\right|\times H\times W\times C+N_{bbx}\times 7)\times 32/8)},$
where $\left|M\right|$ represents the feature retention ratio, $N_{bbx}$ denotes the number of bounding boxes per agent, and 7 is used to describe the coordinates, dimensions, and angle information of each bounding box. Each value is represented using the float32 data type, with division by 8 converting the result to bytes.
Implementation details. We implement our model on the PyTorch toolbox [27] and train it on NVIDIA GeForce RTX 4090 GPUs. We use the Adam optimizer and adopt batch sizes of 3, 5, and 3 for the OPV2V, DAIR-V2X, and V2XSet datasets, respectively, with 60 epochs for each. All models use the PointPillars [15] backbone to extract features from the raw point clouds. The confidence ratio selected for the Confidence-based Filter Generation Module is 70% for OPV2V, 60% for DAIR-V2X, and 70% for V2XSet, and the surrounding grid size chosen for the Multi-scale Offset-aware Attention is 3×3. To closely simulate real traffic conditions, we introduce localization and heading errors with standard deviations of 0.2 $m$ and 0.2°, respectively, sampled from a Gaussian distribution into the initial state of each agent. The communication range is set to 70 $m$. Agents beyond the communication range do not participate in the collaborative process.
4.2 Quantitative Evaluation
Detection Performance Evaluation. Table 1 presents the comparison of the perception performance of our proposed mmCooper model and various baseline models on the OPV2V, DAIR-V2X and V2XSet datasets with a time delay of 100 $ms$, localization errors of 0.2 $m$, and heading errors of 0.2°. The baseline models included in the comparison are described as follows. The No Fusion model relies solely on local observations for 3D object detection. Late Fusion shares the perception results of individual predictions among agents. Intermediate Fusion allows agents to share complete features extracted from raw point clouds. The above three models are all variants based on the PointPillar[15] point cloud detector. Additionally, we consider state-of-the-art (SOTA) models, including When2com [23], DisoNet [18], Where2comm [11], V2XViT [42] and ERMVP [50]. As shown in the table, our proposed mmCooper outperforms previous methods on both simulated and real-world datasets, demonstrating the effectiveness of our approach. Compared to existing SOTA models, mmCooper outperforms the second-best baseline on the OPV2V, DAIR-V2X, and V2XSet datasets by 7.29%/0.59%, 1.31%/0.91% and 2.09%/0.77% in AP@0.7/0.5, respectively. Notably, mmCooper significantly reduces communication costs, achieving levels comparable to late fusion on both datasets. In comparison to other models, the communication volume required to achieve optimal performance is 9153/156/18305 times that of our method. On the OPV2V dataset, our method even achieves lower communication volume than late fusion. A possible explanation is that our Confidence-based Filter Generation Module first removes low-quality features and bounding boxes, then dynamically allocates intermediate-stage features and late-stage bounding boxes based on this. The dynamic allocation process prioritizes using bounding boxes to convey information, preserving a few key features while eliminating a large amount of background information.
Noise Type Localization Errors(m) Transmission Delay(ms)
Datasets OPV2V DAIR-V2X OPV2V DAIR-V2X
Noise Level 0.0 0.2 0.4 0.0 0.2 0.4 0 200 400 0 200 400
No Fusion 48.66 48.66 48.66 43.57 43.57 43.57 48.66 48.66 48.66 43.57 43.57 43.57
When2com [23] 70.82 64.35 59.98 39.50 36.62 35.03 70.82 43.96 36.55 39.50 36.83 33.72
DiscoNet [18] 76.05 73.59 65.85 46.94 46.03 44.96 76.05 60.08 50.33 46.94 44.03 41.57
Where2comm [11] 78.47 75.45 69.77 52.34 49.34 46.96 78.47 65.23 53.33 52.34 47.76 45.25
V2XViT [42] 77.83 75.21 66.72 46.12 43.81 42.53 77.83 66.08 50.33 46.12 45.49 43.99
ERMVP [50] 80.55 76.19 72.78 53.27 48.66 45.97 80.55 68.89 68.31 53.27 49.68 48.60
Ours 86.41 82.53 76.80 56.06 51.52 47.66 86.41 77.44 75.26 56.06 50.96 50.81
Table 2: Results for different localization errors ($m$) and transmission delay ($ms$) on the OPV2V and DAIR-V2X dataset evaluated using AP@0.7. Bold numbers indicate the best results, while underlined numbers represent the second-best results.
Robustness to Localization Errors and Transmission Delay. We verified the robustness of the agents during collaborative processes in the presence of localization errors and transmission delay on the OPV2V and DAIR-V2X datasets. Following the noise settings in [42], localization noise was sampled from a Gaussian distribution with a mean of 0 $m$ and a standard deviation $\sigma\in\{0,0.2,0.4\}m$. As shown in table 2, as the standard deviation of the localization noise gradually increases, the detection performance of the models correspondingly decreases. On the DAIR-V2X dataset, when2com even performs worse than the No Fusion method when the localization error exceeds 0.2 $m$. Our method outperforms previous SOTA models under all localization noise settings, effectively demonstrating its robustness to varying levels of localization error. Transmission delay can lead to misalignment of features and proposals, thereby impacting detection performance. We evaluated the models on the OPV2V and DAIR-V2X datasets with time delays $\tau\in\{0,200,400\}ms$ set to 0 $ms$, 200 $ms$, and 400 $ms$. With increasing time delay, the detection performance of all methods gradually degrades due to misleading features and proposals. However, mmCooper achieves higher Average Precision (AP) than current SOTA models on both datasets. This robustness to localization errors and transmission delay demonstrates that considering spatial continuity aligns features effectively, and intermediate-stage features provide significant guidance for subsequent bounding boxes.
4.3 Ablation Study
GSP MOA FBCFormer LF IF AP@0.7/0.5()($\uparrow$)( ↑ )
OPV2V DAIR-V2X
78.11/88.93 48.27/65.12
Importance Of Core Components
73.46/88.84 47.66/63.63
72.60/88.51 47.27/63.49
76.77/88.56 44.31/59.68
Results of Single-stage Fusion
76.83/88.65 44.11/59.92
66.80/78.62 46.21/57.02
Table 3: Ablation study results of different designs in mmCooper on the OPV2V and DAIR-V2X datasets. GSP: Gumbel Softmax-based Partitioner; MOA: Multi-scale Offset-aware Attention; FBCFormer: Feature-guided Bounding Boxes Calibration Transformer; LF: Late-stage Fusion; IF: Intermediate-stage Fusion.
We conducted comprehensive ablation studies on the OPV2V and DAIR-V2X datasets to demonstrate the importance of different designs, with the ablation results shown in Table 3.
Importance of Core Components. All our core components contribute to performance improvement. Removing the Gumbel Softmax-based Partition (GSP) from the Confidence-based Filter Generation Module leads to a performance drop. This indicates that low-quality information exists in the intermediate-stage features and the late-stage bounding boxes, potentially due to irrelevant background or noise. Introducing adaptive and dynamic allocation for intermediate-stage features and late-stage bounding boxes effectively filters out irrelevant information. The absence of the intermediate-stage multi-scale Offset-aware Attention (MOA) module and the late-stage Feature-guided Bounding Boxes Calibration Transformer (FBCFormer) module also leads to a loss of accuracy, demonstrating that incorporating spatial neighborhood information and cross-stage calibration modules effectively addresses potential misalignment and calibration noise.
Results of Single-stage Fusion. We entirely removed the Late Fusion (LF) and the Intermediate Fusion (IF) separately to degrade our model to single-stage fusion. The experimental results indicate that removing either module leads to a performance decline, underscoring the effectiveness of the proposed multi-stage cooperation approach.
4.4 Qualitative Evaluation
Visualization of Detection Results. To visually display the detection results of our proposed mmCooper and the other baseline models, Figure 6 shows the visualization results for two scenarios in the DAIR-V2X dataset. Compared to SOTAs, our mmCooper achieves high-precision 3D object detection in these challenging scenarios, almost correctly predicting all ground truth. Specifically, the advantages of our method stem from two aspects: 1) The proposed dynamic adaptive allocation strategy enables dynamic partitioning of intermediate-stage features and late-stage bounding boxes, effectively filtering out irrelevant and noisy information. 2) Our Multi-scale Offset-aware Attention Module and BBox Filtering & Calibration Module effectively address misalignment and calibration noise.
Figure 6: Visualization comparison of detection results on the DAIR-V2X dataset. Green and red boxes represent the ground truth and the model-predicted bounding boxes, respectively.
5 Conclusion
In this paper, we have proposed mmCooper, a novel multi-agent, multi-stage, communication-efficient, and collaboration-robust cooperative perception framework. mmCooper is the first framework to achieve adaptive fusion of complementary data across stages for cooperative perception. We introduce the Confidence-based Filter Generation Module to enable dynamic partitioning of intermediate features and late-stage bounding boxes, balancing communication load and perception performance. Additionally, the Multi-scale Offset-aware Fusion module and BBox Filtering & Calibration module are incorporated to address potential misalignment and calibration noise among agents. mmCooper outperforms the second-best state-of-the-art models on the OPV2V, DAIR-V2X, and V2Xset datasets, achieving improvements of 7.29%, 1.31%, and 2.09% in AP@0.7, respectively, while reducing communication volume by factors of 9153, 156 and 18305.
Supplementary Material
1 Overview
The supplementary material is organized into the following sections:
  • Sec. 2: The System Pipeline of mmCooper
  • Sec. 3: Additional Experimental Results on OPV2V and DAIR-V2X Datasets
    • Sec. 3.1: Implementation Details
    • Sec. 3.2: Supplements on Localization Errors
    • Sec. 3.3: Supplements on Transmission Delays
    • Sec. 3.4: Robustness to Heading Errors
  • Sec. 4: Additional Qualitative Evaluation Results
2 The System Pipeline of mmCooper
The proposed system pipeline of mmCooper is illustrated in Alg.1. Note that the following pipeline is executed in parallel across all agents. For simplicity, we describe the pipeline from the perspective of the ego agent. The ego agent is represented by $i$, while the collaborative agents are denoted by $j$. First, the agent generates BEV features $F_{bev}^{i}$ and $F_{bev}^{j}$ through the Observation Module. In the Broadcasting Module, agents generate initial coarse bounding boxes $B_{coarse}^{i}$ and $B_{coarse}^{j}$, where $b^{i}$ and $b^{j}$ represent the number of bounding boxes predicted by the ego agent and the collaborating agents, respectively. The collaborative agents then package and broadcast the filtered BEV features along with the coarse bounding boxes. Specifically, $f_{gaussian}(\cdot)$ denotes a Gaussian filter, and $\{F_{bev,m}^{j},B_{m}^{j}\}$ represents the filtered features and bounding boxes from the collaborative agents.
Subsequently, in the Intermediate-stage Fusion, the ego agent performs feature fusion using the Multi-scale Offset-aware Attention module. The outputs from the fused features at different scales, after undergoing upsampling, are concatenated to obtain the final fused features. Here, $f_{up2}(\cdot)$ and $f_{up3}(\cdot)$ denote the upsampling operations, while $\{F_{i}^{sc1},F_{i}^{sc2},F_{i}^{sc3}\}$ represent the feature fusion outputs at three different scales.
In the later-stage fusion, the BBox Filtering & Calibration Module is employed to learn the bounding box offsets and scores. Specifically, DBA refers to the Deformable Bounding Boxes Attention Module, while FFN denotes the feed-forward network, $f_{enc,b}(\cdot)$ represents the feature extractor for bounding boxes, $f_{off}(\cdot)$ is the offset mapping function, and $f_{score}(\cdot)$ is the score mapping function. $\varphi(\cdot)$ represents the filtering and calibration through scores and offsets.
Finally, the fused features are input into a detection head to predict the fused bounding box $\mathcal{B}_{fused}$. The final bounding box is composed of the initial prediction, the bounding box predicted from the fused features, and the bounding box obtained after filtering and calibration from collaborative agents. $f_{post}(\cdot)$ represents the combination of all bounding boxes.
Algorithm 1 System Pipeline of the Proposed mmCooper
Define $N=\{1,...,n\}$ as the agent set, $i\in N$ represents the ego agent, while $j\in N$ denotes the collaborative agents. $X_{i/j}$ serves as the input point cloud.
# For collaborative agents.
# Observation Module.
for each agent $j\in N$, do
$X_{j}^{{}^{\prime}}=T_{j->i}\cdot X_{j}$
$F_{bev}^{j}=\psi_{encoder}(X_{j}^{{}^{\prime}})\in\mathbb{R}^{C\times H\times W}$
end for
# Broadcasting Module.
for each $j\in N$, do
$B_{coarse}^{j}=f_{dec}(F_{bev}^{j})\in\mathbb{R}^{b^{j}\times 7}$
$C_{f}^{j},C_{b}^{j}=\phi_{conf}\left(F_{bev}^{j}\right)$
$C_{f}^{j},C_{b}^{j}=f_{gaussian}\left(C_{f}^{j},C_{b}^{j}\right)$
$M_{f,top}^{j},M_{b,top}^{j}=f_{top}\left(C_{f}^{j},C_{b}^{j}\right)$
$M_{f,g}^{j},M_{b,g}^{j}=f_{max}\left(\frac{exp\left((logC_{f/b}^{j}+g_{f})/\tau\right)}{{{\textstyle\sum_{s}}exp\left((logC_{s}^{j}+g_{s})/\tau\right)}}\right)$
$\mathcal{M}_{f/b}^{j}=M_{f/b,g}^{j}\odot M_{f/b,top}^{j}$
$F_{bev,m}^{j}=\mathcal{M}_{f}^{j}\odot F_{bev}^{j}$; $B_{m}^{j}=\mathcal{M}_{b}^{j}\odot B_{coarse}^{j}$
broadcast $\{F_{bev,m}^{j},B_{m}^{j}\}$ to other agent
end for
# For ego agent.
# Observation Module.
$F_{bev}^{i}=\psi_{encoder}(X_{i})\in\mathbb{R}^{C\times H\times W}$
# Broadcasting Module.
$B_{coarse}^{i}=f_{dec}(F_{bev}^{i})\in\mathbb{R}^{b^{i}\times 7}$
# Intermediate-stage Fusion.
Receive $\{F_{bev,m}^{j},B_{m}^{j}\}$ sent by collaborative agents.
Encode the feature set $\{F_{bev}^{i},F_{bev,m}^{j}\}$ into three scales $S=\{sc1,sc2,sc3\}.$
For each $sc\in S$, do
$\mathcal{F}_{i,c}^{sc}=CrossAttn(\sigma(F_{bev,c}^{i,sc}),F_{bev,c}^{j,s\times s,sc},F_{bev,c}^{j,s\times s,sc})$
end for
$\mathcal{F}_{i}=concat(\mathcal{F}_{i}^{sc1},f_{up2}(\mathcal{F}_{i}^{sc2}),f_{up3}(\mathcal{F}_{i}^{sc3}))$
# Late-stage Fusion.
$F_{b}(q)=f_{enc,b}(B_{i,m})$
$DBA(q)=\sum_{\alpha=1}^{A}W_{\alpha}[\sum_{n=1}^{N}\sum_{m=1}^{M}\omega(W_{\beta}F_{b}(q))F_{j}(q+\bigtriangleup q_{m})]$
$F_{DBA}=DBA(q)+\mathcal{F}_{i}$
$F_{DBA}=FFN(F_{DBA})+F_{DBA}$
$off,score=f_{off}(F_{DBA}),f_{score}(F_{DBA})$
$\mathcal{B}_{m}^{j}=\varphi(off,score,B_{m}^{j})$
# Detection Decoders
$\mathcal{B}_{fused}=f_{dec}(\mathcal{F}_{i})$
$\mathcal{B}_{final}^{i}=f_{post}(B_{coarse}^{i},\mathcal{B}_{m}^{j},\mathcal{B}_{fused})$
3 Additional Experimental Results on OPV2V and DAIR-V2X Datasets
In this section, we provide additional experiments to supplement the results on the datasets OPV2V[43] and DAIR-V2X[49] presented in the main paper.
3.1 Implementation Details
On the OPV2V and DAIR-V2X datasets, the dimensions of the voxels encoded by the encoder are $0.4\times 0.4\times 4$. The shape of the shared BEV features among agents is (64, 100, 252) for DAIR-V2X and (64, 100, 352) for OPV2V. The shared bounding boxes among agents are represented by their center coordinates, dimensions (length, width, height), and heading angle. The detection decoder consists of two distinct 1×1 convolutional layers.
3.2 Supplements on Localization Errors
The localization errors on the OPV2V and DAIR-V2X datasets are sampled from a Gaussian distribution with a mean of 0 $m$ and a standard deviation $\sigma\in\{0.0,0.1,0.2,0.3,0.4\}m$. The experimental results, as shown in Figure 7, demonstrate that our proposed mmCooper outperforms the existing state-of-the-art methods[23, 11, 18, 42, 50] across different levels of localization error, highlighting its robustness to such errors.
3.3 Supplements on Transmission Delays
We evaluated the performance of the models on the OPV2V and DAIR-V2X datasets under transmission delays for $\{0,100,200,300,400\}ms$. As shown in Figure 8, our proposed mmCooper consistently outperforms the existing state-of-the-art methods across all transmission delay conditions, demonstrating the superiority of our approach in scenarios with transmission delays.
Figure 7: Robustness to the localization error on the OPV2V and DAIR-V2X datasets.
Figure 8: Robustness to the transmission delay on the OPV2V and DAIR-V2X datasets.
Figure 9: Robustness to the heading error on the OPV2V and DAIR-V2X datasets.
dui
3.4 Robustness to Heading Errors.
Figure 9 demonstrates the performance of our proposed mmCooper method compared to other baseline methods under varying levels (i.e., $\{0.0,0.2,0.4,0.6,0.8\}^{\circ}$) of heading error noise on the OPV2V and DAIR-V2X datasets. As illustrated in Figure 9, the performance of all models decreases as the heading errors increase. However, mmCooper consistently outperforms the current state-of-the-art models across all error levels, highlighting the advantages of our designed Multi-Scale Offset-Aware Fusion Module and BBox Filtering & Calibration Module in enhancing system robustness.
4 Additional Qualitative Evaluation Results
We provide additional qualitative results on DAIR-V2X datasets. Figures 10 present visualizations of different road scenarios. Our proposed mmCooper method can detect almost all ground truths without any false positives. The results demonstrate that our proposed mmCooper achieves outstanding detection performance across various scenarios.
Figure 10: More visualization comparison of detection results on the DAIR-V2X dataset. Green and red boxes represent the ground truth and the model-predicted bounding boxes, respectively.